WHERE MILLIMETER WAVE SPECTRA ARE SENSITIVE TO SMALL ELECTRIC FIELDS: HIGH RYD-BERG STATES OF XENON AND THEIR HYPERFINE STRUCTURES

\underline{MARTIN SCHÄFER}, MATTHIAS RAUNHARDT, and FRÉDÉRIC MERKT, ETH Zürich, Laboratorium für Physikalische Chemie, Wolfgang-Pauli-Strasse 10, 8093 Zürich, Switzerland.

In the range 0–45 cm^{−1} below the ionization limit, the separation between adjacent electronic states (Rydberg states with principal quantum number n > 50) of atoms and molecules is smaller than 2 cm^{−1}. In order to resolve the fine or hyperfine structure of these states, it is necessary to combine high-resolution vacuum ultraviolet (VUV) laser radiation, which is required to access the Rydberg states from the ground state, with millimeter wave radiation.^{a} Such double-resonance experiments have been used to study the hyperfine structure of high Rydberg states of ^{83}Kr ^{b}, H_{2} ^{c}, or D_{2} ^{d}. Millimeter wave transitions (240–350 GHz) between nℓ (52 ≤ n ≤ 64, ℓ ≤ 3) Rydberg states of different xenon isotopes were detected by pulsed field ionization followed by mass-selective detection of the cations. Because of the high polarizability of high-n Rydberg states (∝ n^{7}, ∼ 10^{4} MHz cm^{2} V^{−2} for n ≈ 50), it is necessary to reduce the electric stray fields to values of the order of mV/cm (or less) in order to minimize the (quadratic) Stark shift of the millimeter wave transitions. Some p and d Rydberg states of Xe are nearly degenerate and efficiently mixed by small stray fields, making it possible to observe transitions forbidden by the Δℓ = ±1 selection rule or transitions exhibiting a linear Stark effect, which is typical for the degenerate high-ℓ Rydberg states. Multichannel quantum defect theory (MQDT) was used to analyze the millimeter wave data and to determine the hyperfine structures of the ^{2}P_{3/2} ground electronic states of ^{129}Xe^{+} and ^{131}Xe^{+}.

aC. Fabre, P. Goy, S. Haroche, J. Phys. B: Atom. Mol. Phys. 10, L183–189 (1977). F. Merkt, A. Osterwalder, Int. Rev. Phys.

Chem. 21, 385–403 (2002). M. Schäfer, M. Andrist, H. Schmutz, F. Lewen, G. Winnewisser, F. Merkt, J. Phys. B: At. Mol. Opt.

Phys. 39, 831–845 (2006).

bM. Schäfer, F. Merkt, Phys. Rev. A, 74, 062506 (2006).

cA. Osterwalder, A. Wüest, F. Merkt, Ch. Jungen, J. Chem. Phys., 121, 11810–11838 (2004).

dH. A. Cruse, Ch. Jungen, F. Merkt, Phys. Rev. A 77, 04502 (2008).